Asim is now 3times as old as irfan . After 10 years. Asim will be twice as old as irfan .Asim age at this time is.
Found 4 solutions by saw, MathTherapy, Alan3354, ikleyn:
Answer by saw(34)

I = x
A = 3x
After 10 years
I = x+10
A = 2(x+10) = 2x+20
(x+3x)+10 = (x+10+2x+20)
4x+10 = 3x+30
x = 20
3x = 60
Check-----
(x+3x)+10 = (x+10+2x+20)
90 = 90


Answer by MathTherapy(10552)


Asim is now 3times as old as irfan . After 10 years. Asim will be twice as old as irfan .Asim age at this time is.
The other person is WRONG!
Asim is NOT 60 - FAR from it, and Irfan is NOT 20. Again, far from it!

Answer by Alan3354(69443)

Asim is now 3 times as old as irfan. After 10 years, Asim will be twice as old as irfan. Asim age at this time is.
====================
A = 3I
A+10 = 2(I+10) = 2I+20
3I+10 = 2I+20
I = 10
A = 30

Answer by ikleyn(52781)

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.